# Improving Neutral Point of View Text Generation through Parameter-Efficient Reinforcement Learning and a Small-Scale High-Quality Dataset

### v1

@@ Abstract.
This paper describes the construction of a dataset and the evaluation of training methods to improve generative large language models’ (LLMs) ability to answer queries on sensitive topics with a Neutral Point of View (NPOV), i.e., to provide significantly more informative, diverse and impartial answers. The dataset, SHQ-NPOV, comprises 300 high-quality, human-written quadruplets: a query on a sensitive topic, an answer, an NPOV rating, and a set of links to source texts elaborating the various points of view. The first key contribution of this paper is a new methodology to create such datasets through iterative rounds of human peer-critique and annotator training, which we release alongside the dataset. The second key contribution is the identification of a highly effective training regime for parameter-efficient reinforcement learning (PE-RL) to improve NPOV generation. We compare and extensively evaluate PE-RL and multiple baselines—including LoRA finetuning (a strong baseline), SFT and RLHF. PE-RL not only improves on overall NPOV quality compared to the strongest baseline ($97.06\%\rightarrow 99.08\%$), but also scores much higher on features linguists identify as key to separating good answers from the best answers ($60.25\%\rightarrow 85.21\%$ for presence of supportive details, $68.74\%\rightarrow 91.43\%$ for absence of oversimplification). A qualitative analysis corroborates this. Finally, our evaluation finds no statistical differences between results on topics that appear in the training dataset and those on separated evaluation topics, which provides strong evidence that our approach to training PE-RL exhibits very effective out of topic generalization.
@@
Neutral Point of View, Low-data, Parameter-efficient methods, Parameter-efficient Reinforcement Learning, PE-RL, LoRA
ccs: Computing methodologies Natural language generation
1. Introduction
@@
Figure 1. Pipeline to create Neutral Point of View (NPOV) answers to queries on sensitive topics.
@@
A long standing goal of AI question-answering systems has been to provide multiple perspectives to controversial queries (Cardie et al., 2003; Ku and Chen, 2007; Sun et al., 2017; Chen et al., 2022, 2019). Applications aimed at summarizing diverse opinions range from helping people make more informed choices on product purchases (Hu and Liu, 2004) to predicting elections(Tumasjan et al., 2010). With the development of chatbots and LLMs playing an increasingly significant role in search and question answering, several authors such as Metzler et al. (2021) have argued that there is growing importance to provide diverse, unbiased points of view. Doing this well may help with the challenges of polarization and misinformation (Costello et al., 2024), and failure has the risk of exacerbating social tensions. This motivates the question of how to efficiently and effectively control LLMs to provide high quality, multi-perspective responses to controversial topics.
@@
In this paper, we addresses this challenge by first describing a methodology we developed to create a small high quality dataset of responses to controversial topics with varying degrees of neutrality; and second by investigating training regimes to identify one that not only learns to generate high quality, neutral, multi-perspective responses, but also successfully generalises to out of domain topics.
@@
The dataset we use as a case study for our method consists of 300 examples with high inter-annotator agreement, and is based on adapting Wikipedia’s Neutral Point of View (NPOV) policy. It includes a diverse spectrum of responses exhibiting varying degrees of bias, informativeness, clarity of the arguments, and over-simplification, among other features. The most effective training regime we identified is parameter-efficient reinforcement learning (PE-RL) (Sidahmed et al., 2024). While this training algorithm has received very little attention, it both improves on overall NPOV quality compared to the strongest baseline ($97.06\%\rightarrow 99.08\%$), and also scores much higher on features which linguists identify as being key to separating good answers from the best answers ($60.25\%\rightarrow 85.21\%$), for presence of supportive details ($68.74\%\rightarrow 91.43\%$), for absence of oversimplification. We compare this methods to strong baselines, including LoRA SFT (parameter-efficient Supervised FineTuning), best-of-10, traditional RLHF (Reinforcement Learning from Human Feedback) and traditional SFT (Supervised FineTuning). Perhaps the most surprising result of our evaluation is that PE-RL, unlike the non-parameter-efficient methods (including RLHF), exhibits no perceptible over-fitting, despite the small-data regime during training. When evaluated on 300 additional questions spanning 30 different topics sourced from a different set of sites, the PE-RL based LLM generator exhibits equally good responses.
@@
1.1. Related work
Perhaps the most closely related work is Wikipedia’s NPOV policy; see Section 2.2 for in-depth similarities and differences with our work. In Wikipedia, it is possible to flag articles as not NPOV. Pavalanathan et al. (2018) shows NPOV flagging and subsequent edits help reduce biased language, but has little success training individual editors. Using the edits, Recasens et al. (2013) identifies words that make sentences not neutral. They also create the NPOV corpus, a dataset of edits meant to remove bias. This dataset can be used to rewrite sentences in a more NPOV way; in contrast, our dataset directly provides answers, accompanied with an NPOV score. Pryzant et al. (2020) tackles a similar task, and releases another dataset of pairs of biased/neutralized sentences. Other datasets of interest for our work can be found in (Slonim et al., 2021; Bar-Haim et al., 2020; Sznajder et al., 2019), which include large corpora of arguments covering sensitive topics (as a side-product of creating debating systems). Finally, Sun et al. (2023) creates a corpus of controversial query-answers for evaluating models, subsampled from the well-known Quora QuestionPair dataset (Wang et al., 2017); in contrast, our dataset contains only original content.
Neutral Point of view text generation is closely related to the field of multiple perspectives answering–a common way of increasing the diversity of LLMs’ outputs (Patankar et al., 2018; Lahoti et al., 2023). Metzler et al. (2021) argues that that LLMs’ responses should include diverse perspectives with some degree of impartiality: ”Generated responses should represent a range of diverse perspectives but should not be polarizing. For example, for queries about controversial topics, both sides of the topic should be covered in a fair and balanced way.” Chen et al. (2022) comments on the difficulty of the multiple-perspective answering task, and state that debate-worthy, controversial questions are challenging to design information retrieval systems for. Similar to (Metzler et al., 2021), they also state that ”a retrieval system should recognize the semantic difference of responses in cross-document settings, and in turn organize and deliver a set of documents from diverse perspectives.”.
Providing multi-perspective answers is a significant branch of work in the field of information retrieval  (Cardie et al., 2003; Metzler et al., 2021; Chen et al., 2022). The study by Chang et al. (2024) delves into managing controversial discussions within LLM-based chatbots by adhering to Wikipedia’s Neutral Point of View (NPOV) principle. It introduces a retrieval-augmented generation framework that leverages multiple perspectives retrieved from a knowledge base. The study identifies and addresses common LLM failures such as hallucination and coverage errors, proposing three detection methods based on word overlap, salience, and LLM-based classifiers. In a related field, to increase the controllability of LLMs on controversial topics, Li et al. (2024) makes LLMs debate between themselves to produce diverse different perspectives, and Chen et al. (2019) proposes answering with multiple perspectives backed by facts, along a spectrum. The Bing search engine implemented a specialised view for multiple-perspective answering in 2018, arguing that it helps address echo chamber effects and search-based confirmation bias  [^1] https://blogs.bing.com/search-quality-insights/february-2018/toward-a-more-intelligent-search-bing-multi-perspective-answers.
Finally, by tackling controversial topics identified by outside sources (see Section 2.2 for details), we implicitly present perspectives that are backed by a significant portion of the population. However, identifying controversial topics is a research problem by itself, tackled for instance by (Chen et al., 2023; Sznajder et al., 2019; Kittur et al., 2007; Popescu and Pennacchiotti, 2010).
2. Dataset creation and annotation process
2.1. Dataset overview
@@
Table 1. Examples of (Query, Answer, Score) for a given topic: ”lowering the drinking age in the US”. The score (0 to 5) quantifies the NPOVness of the answer, regardless of whether or not the query is adversarial. An answer is classified as NPOV if its score is $\geq 3$.
The dataset we create—the SHQ-NPOV dataset [^2] https://docs.google.com/spreadsheets/d/11grN4EoCuj6DLuTtykyJek8BkuwLXI5toHUu9lV0iE8/—is a list of “controversial” human-written queries and answers from varying “points of view”: both queries and answer are sometimes neutral, sometimes taking a side (queries and answers can take opposite sides). A group of 4 expert writers from diverse demographics (see Appendix D for details) developed (wrote, peer-critiqued, updated etc.) this dataset used for training models. It contains 300 entries (see Table 1 for 4 entries on the same topic), each comprising:
@@
The dataset also contains 1190 additional queries without answers (see Table 3). We evaluate all our models on the responses to these prompts. Those queries comprise 10 templates covering the 89 topics of the SHQ-NPOV dataset (which we call in-distribution topics), and 30 additional topics (which we call out-of-distribution topics).
[0, 1] (1, 2] (2, 3) [3, 4) [4, 5]
perspective or not NPOV informative or style could be well-
insulting language subtly not NPOV improved written
Examples in bucket(of dataset) 88 (29.33%) 20 (6.67%) 20 (6.67%) 40 (13.33%) 132 (44%)
NPOV examples (of dataset) Not NPOV: 128 (42.67%) NPOV: 172 (57.33%)
Table 2. Dataset characteristics.
2.2. Defining the NPOV task
Our task is similar to the Neutral Point of View response task from (Chang et al., 2024), which itself was inspired by Wikipedia’s policy. We cite here verbatim the relevant parts of the first section of Wikipedia’s page on Neutral Point of View [^3] Emphasis and colors are from the webpage., which all of our writers read and attempt to follow when developing their examples:
Avoid stating opinions as facts.:
Avoid stating seriously contested assertions as facts.:
Avoid stating facts as opinions.:
Prefer nonjudgmental language.:
Indicate the relative prominence of opposing views.:
One fundamental goal of this work is to combat polarization by developing a deeper understanding of different points of view and empathy for the people who express them. As such, we strive to ensure that people’s views are represented; this can be seen as a form of reflective listening, which has been shown to decrease defensiveness and increase willingness to hear different perspectives (Dalmar, 2007). As a consequence, we do not assume a hierarchy between scientific arguments and emotional or spiritual arguments; this is in stark contrast with the goal of Wikipedia or any encyclopedia. In our definition of NPOV, by design, peer-reviewed quantitative studies can be presented on equal ground with widespread moral positions.
@@ 2.2. Defining the NPOV task
2.3. Methodology for writing examples
To achieve the task detailed above, we take the following steps:
Gather a diverse team: On top of any conscious biases a person can have, they also bring unconscious biases depending on their identity and their cultural background, which clashes with the goal of creating NPOV answers. While it is arguably impossible for a single person to write impartial answers, we try to mitigate this effect by involving a group of people at each step of the writing. We gather a group of 4 writers of different gender, race, sexual orientation, religion and country of residence (see Appendix D for details) [^4] While the group of 4 people was diverse along multiple axis, they were similar in age and income bracket status. 4 people also cannot possibly represent all the diversity of existing races, gender identities, countries and sexual orientation (among others). We see this dataset as a first step in the NPOV direction, and we hope it will one day become part of a more diverse and richer dataset.. Each of these writers writes examples and peer-critiques other writers’ examples according to the NPOV policy.
Figure 2. Life of an example.
Writing, peer-critiques and revisions:
(see Figure 2). Each writer has their own document, accessible by everyone with the option to add comments. This documnent comprises 6 sections: ”In progress”, ”New example”, ”In revision”, ”Revised” and ”Approved”, ”Added to dataset”. They work on their draft in the ”In progress” section. When an example is finished, it is moved to the ”New example” section. Here, peers (usually 2 of them) read it and either deem it ready—in which case they move it to the ”Approved” section— or leave both high-level comments (e.g. if important arguments are missing) and detailed comments (e.g. if a choice of words is not neutral)—in which case the example is moved to ”In revision”. In the latter case, the initial writer revises the example based on the comments, and moves the example to the ”Revised” section when done. This example is then moved to the ”Approved” section if ready, or moved back to the ”In revision” section as many times as needed until it gains the peers’ approval. Finally, the writer moves the example to ”Added to dataset” when they have copied it to an external document gathering all the writers’ examples. Note that the ”New example” and ”Revised” sections are similar, in the sense that examples are waiting for peer critique. However, we found that new examples usually need more careful examination as well as more comments than revised answers, which usually only require checking whether previous comments were addressed. Separating the 2 sections therefore increased peer critique efficiency.
Writing non-NPOV answers Writers found writing blatantly non-NPOV examples cathartic and enjoyable, reporting increased motivation for writing NPOV examples afterward. Writers were instructed to alternate between non-NPOV answers that they agree with and non-NPOV answers they disagree with. In both case, writing non-NPOV examples was much faster than writing NPOV examples. The non-NPOV examples are part of the dataset, as some methods utilize negative examples.
2.4. Crowd-annotation procedure
Figure 3. Distribution of L1 distance between annotators’ score and NPOV score. 75% of answers have less than 0.50 difference with the NPOV score, and 90% of answers less than 1.
@@ 2.4. Crowd-annotation procedure
To validate our annotation task, we hired a team of 10 external crowd-annotators based in the US and recruited 4 internal experts, for a total of 14 annotators. 2 experts first annotated the whole dataset, using a score from 0 to 5 to capture ”NPOVness” (see Table 2 for score meaning). After reading the Wikipedia definition for NPOV, crowd-annotators are then split into 2 groups, and trained for the task on half the dataset. They are then asked to rate the NPOVness of an answer for the other half. At least 3 annotators rate each query-answer pair across all experiments, for a maximum of 5. The final NPOV score is the average of the 3–5 annotators’ scores. See Figure 6 in the Appendix.
@@
The annotators have high agreement, as seen in the metrics below. When we compute a metric with respect to the NPOV score (defined as the average of the annotator’s score), the metric is unfairly high, as every annotator contribute to this score. Similarly, when we compute a metric with respect to the score of a single expert, the metric is unfairly low. We can therefore make the following claims:
  • The overall annotator average accuracy is within $[91.07\%,91.15\%]$.
  • The overall annotator average AUC is within $[96.53\%,97.59\%]$.
@@
The last point is of particular interest. Most annotators’ score is less than 0.5 away from the real score, which is very precise for a 0–5 score. We plot the exact distribution of L1 distances:
2.5. 1190 extra queries for evaluation
Question templates Opinionated? Personal?
Can you explain to me why there is a polemic about [topic]? Neutral No
Tell me more about [topic] Neutral No
How do you feel about [topic]? Neutral Yes
What do you think of [topic]? Neutral Yes
I oppose [topic]. Can you give me some arguments against it? Con No
Everyone should back up [topic], right? Pro No
Can you give me some argument against [topic]? Con No
I support [topic], what about you? Pro Yes
Are you against [topic]? Con Yes
Table 3. Templates of the queries for evaluation.
The SHQ-NPOV dataset consists of 300 query-answer pairs covering 89 topics, which we call in-distribution topics (id topics). On top of it, we also build a query dataset for evaluation using both these topics and 30 out-of-distribution ones (ood topics), making a total of 119 that we upsample using the 10 templates from Table 3, giving a total of 1190 evaluation queries. Note that there is no overlap between the extra queries and the queries in the SHQ-NPOV dataset.
@@
3. Experimental setup
3.1. Methods and baselines
We now describe the different methods we compare for generating NPOV answers. Note that all the methods are used to tune the same base model. We compare the results of PE-RL (Parameter-Efficient Reinforcement Learning) to the following baselines: base model, LoRA (parameter-efficient Supervised Fine Tuning), SFT (”traditional” Supervised Fine Tuning), RLHF (”traditional” Reinforcement Learning with Human Feedback) and best-of-10. For each of these techniques, we also consider its variant ”+ preamble”, which consists of adding the preamble in Appendix C.1 before the query.
3.1.1. PE-RL
Following the method introduced in (Sidahmed et al., 2024), we train 2 LoRA adapters, one for the reward model and one for the policy. We first train the reward model by finetuning LoRA (lora) on the base model. We use a rank-4 LoRA adapter (about 4M trainable parameters), a learning rate of $0.0005$ for $8,000$ steps to train on all the 300 rated examples. We then optimize the RL loop and update the parameters of the rank-4 LoRA adapter for the policy using the trained reward model for $2,000$ steps (with 1,000 warm-up steps), using a value learning rate of $0.0001$, policy learning rate of $0.00001$. We perform a parameters search to pick the best model. The LoRA adapter for the policy is initialized at 0 at the begininng of the RL loop.
3.1.2. PE-RL + LoRA SFT
Instead of initializing the LoRA adapter for the policy with a null checkpoint, we perform LoRA SFT on the 132 great (query, answer) pairs, and initialize the RL training with this trained checkpoint. Note that the 132 examples are part of the 300 used for training the reward model, so this method does not use additional data.
3.1.3. PaLM 2
We report the generations from the base model, PaLM 2.
3.1.4. LoRA SFT
We fine-tuned the base model with a rank-4 LoRA (Hu et al., 2021) adapter (about 6M trainable parameters) using only examples rated as good or above ($N=132$), a batch size of $64$, a learning rate of $0.0001$, and dropout probability of $0.1$ for $800$ steps. The set of hyper-parameters is selected from a grid search of LoRA ranks in $\{1,4,8,16\}$, learning rate in $\{0.0004,0.005,0.001,0.05\}$, and dropout probability in $\{0,0.05,0.1,0.2\}$.
3.1.5. SFT
We fine-tuned the base LLM exclusively on examples with ratings of good or higher ($N=132$). The model was trained for $1000$ steps using a batch size of $64$, a learning rate of $0.0001$, and a dropout rate of $0.1$. These hyper-parameters were chosen based on a grid search over learning rates $\{0.0001,0.001,0.005,0.05\}$ and dropout rates $\{0,0.05,0.1,0.2\}$.
3.1.6. RLHF

### v2

# Improving Neutral Point-of-View Generation with Data- and Parameter-Efficient RL

@@ Abstract
The paper shows that parameter-efficient reinforcement learning (PE-RL) is a highly effective training regime to improve large language models’ (LLMs) ability to answer queries on sensitive topics with a Neutral Point of View (NPOV), i.e. to provide significantly more informative, diverse and impartial answers. This is shown by evaluating PE-RL and multiple strong baselines—including LoRA finetuning (strongest baseline), SFT and RLHF. PE-RL not only improves on overall NPOV quality compared to the strongest baseline ($97.06\%\rightarrow 99.08\%$), but also scores much higher on features linguists identify as key to separating sufficient answers from “great” answers ($60.25\%\rightarrow 85.21\%$ for presence of supportive details, $68.74\%\rightarrow 91.43\%$ for absence of oversimplification). A qualitative analysis corroborates this. Moreover, our evaluation also finds a key property of PE-RL for this task: unlike methods that update all parameters, it generalises out of topic. Finally, to enable further studies we also release the dataset, SHQ-NPOV, and provide a methodology to create such datasets through iterative rounds of human peer-critique and annotator training.
@@
Improving Neutral Point-of-View Generation with Data- and Parameter-Efficient RL
Jessica Hoffmann1, Christiane Ahlheim2, Zac Yu2, Aria Walfrand2, Jarvis Jin1, Marie Tano2, Ahmad Beirami1, Erin van Liemt2, Nithum Thain1, Hakim Sidahmed1, Lucas Dixon1 1Google DeepMind, 2Google Correspondence: jhoffmann@google.com
1 Introduction
@@
Figure 1: Pipeline to create Neutral Point of View (NPOV) answers to queries on sensitive topics.
@@
A long standing goal of AI question-answering systems has been to provide multiple perspectives to controversial queries Cardie et al. (2003); Ku and Chen (2007); Sun et al. (2017); Chen et al. (2022, 2019). Applications aimed at summarizing diverse opinions range from helping people make more informed choices on product purchases Hu and Liu (2004) to predicting electionsTumasjan et al. (2010). With the development of chatbots and LLMs playing an increasingly significant role in search and question answering, several authors such as Metzler et al. (2021) have argued that there is growing importance to provide diverse, unbiased points of view. Doing this well may help with the challenges of polarization and misinformation Costello et al. (2024), and failure has the risk of exacerbating social tensions. This motivates the question of how to efficiently and effectively control LLMs to provide high quality, multi-perspective responses to controversial topics.
@@
In this paper, we addresses this challenge by first describing a methodology we developed to create a small high quality dataset of responses to controversial topics with varying degrees of neutrality; and second by investigating training regimes to identify one that not only learns to generate high quality, neutral, multi-perspective responses, but also successfully generalises to out of domain topics. The resulting dataset we use as a case study for our method consists of 300 examples with high inter-annotator agreement, and is based on adapting Wikipedia’s Neutral Point of View (NPOV) policy. It includes a diverse spectrum of responses exhibiting varying degrees of bias, informativeness, clarity of the arguments, and over-simplification, among other features.
@@
The key contribution is the identification of parameter-efficient reinforcement learning (PE-RL) Sidahmed et al. (2024) as a remarkable effective, and understudied paradigm for low-data controlled generation. PE-RL both improves on overall NPOV quality compared to the strongest baseline ($97.06\%\rightarrow 99.08\%$), and also scores much higher on features which linguists identify as being key to separating sufficient answers from great answers ($60.25\%\rightarrow 85.21\%$), for presence of supportive details ($68.74\%\rightarrow 91.43\%$), and for absence of oversimplification. We compare this methods to strong baselines, including LoRA SFT (parameter-efficient Supervised FineTuning), best-of-10, traditional RLHF (Reinforcement Learning from Human Feedback) and traditional SFT (Supervised FineTuning). Perhaps the most surprising result of our evaluation is that PE-RL, unlike the non-parameter-efficient methods (including RLHF), exhibits no perceptible over-fitting, despite the small-data regime during training. When evaluated on 300 additional questions spanning 30 different topics sourced from a different set of sites, the PE-RL based LLM generator exhibits equally good responses.
@@
Table 1: Examples of (Query, Answer, Score) for a given topic: "lowering the drinking age in the US". The score (0 to 5) quantifies the NPOVness of the answer, regardless of whether or not the query is adversarial. An answer is classified as NPOV if its score is $\geq 3$.
2 Related work
Perhaps the most closely related work is Wikipedia’s NPOV policy; see §3.2 for in-depth similarities and differences with our work. In Wikipedia, it is possible to flag articles as not NPOV. Pavalanathan et al. (2018) shows NPOV flagging and subsequent edits help reduce biased language, but has little success training individual editors. Using the edits, Recasens et al. (2013) identifies words that make sentences not neutral. They also create the NPOV corpus, a dataset of edits meant to remove bias. This dataset can be used to rewrite sentences in a more NPOV way; in contrast, our dataset directly provides answers, accompanied with an NPOV score. Pryzant et al. (2020) tackles a similar task, and releases another dataset of pairs of biased/neutralized sentences. Other datasets of interest for our work can be found in Slonim et al. (2021); Bar-Haim et al. (2020); Sznajder et al. (2019), which include large corpora of arguments covering sensitive topics (as a side-product of creating debating systems). Finally, Sun et al. (2023) creates a corpus of controversial query-answers for evaluating models, subsampled from the well-known Quora QuestionPair dataset Wang et al. (2017); in contrast, our dataset contains only original content.
Neutral Point of view text generation is closely related to the field of multiple perspectives answering–a common way of increasing the diversity of LLMs’ outputs Patankar et al. (2018); Lahoti et al. (2023). Metzler et al. (2021) argues that that LLMs’ responses should include diverse perspectives with some degree of impartiality: "Generated responses should represent a range of diverse perspectives but should not be polarizing. For example, for queries about controversial topics, both sides of the topic should be covered in a fair and balanced way." Chen et al. (2022) comments on the difficulty of the multiple-perspective answering task, and state that debate-worthy, controversial questions are challenging to design information retrieval systems for. Similar to Metzler et al. (2021), they also state that "a retrieval system should recognize the semantic difference of responses in cross-document settings, and in turn organize and deliver a set of documents from diverse perspectives.".
Providing multi-perspective answers is a significant branch of work in the field of information retrieval  Cardie et al. (2003); Metzler et al. (2021); Chen et al. (2022). The study by Chang et al. (2024) delves into managing controversial discussions within LLM-based chatbots by adhering to Wikipedia’s Neutral Point of View (NPOV) principle. It introduces a retrieval-augmented generation framework that leverages multiple perspectives retrieved from a knowledge base. The study identifies and addresses common LLM failures such as hallucination and coverage errors, proposing three detection methods based on word overlap, salience, and LLM-based classifiers. In a related field, to increase the controllability of LLMs on controversial topics, Li et al. (2024) makes LLMs debate between themselves to produce diverse different perspectives, and Chen et al. (2019) proposes answering with multiple perspectives backed by facts, along a spectrum. The Bing search engine implemented a specialised view for multiple-perspective answering in 2018, arguing that it helps address echo chamber effects and search-based confirmation bias  [^1] See here..
Finally, by tackling controversial topics identified by outside sources (see §3.2 for details), we implicitly present perspectives that are backed by a significant portion of the population. However, identifying controversial topics is a research problem by itself, tackled for instance by Chen et al. (2023); Sznajder et al. (2019); Kittur et al. (2007); Popescu and Pennacchiotti (2010).
3 Dataset creation and annotations
3.1 Dataset overview
The dataset we create—the SHQ-NPOV dataset —is a list of “controversial” human-written queries and answers from varying “points of view”: both queries and answer are sometimes neutral, sometimes taking a side (queries and answers can take opposite sides). A group of 4 expert writers from diverse demographics (see Appendix I for details) developed this dataset for use in training. It contains 300 entries (see Table 1 for 4 entries on the same topic), each comprising:
@@ 3.1 Dataset overview
$57\%$ of the dataset is labeled NPOV. Precise distribution by score can be found in Table 5 in the Appendix. The dataset also contains 1190 additional queries without answers (see Table 6). We evaluate all our models on the responses to these prompts. Those queries comprise 10 templates covering the 89 topics of the SHQ-NPOV dataset (which we call in-distribution topics), and 30 additional topics (which we call out-of-distribution topics).
3.2 Defining the NPOV task
Our task is similar to the Neutral Point of View response task from Chang et al. (2024), which itself was inspired by Wikipedia’s policy. The task is to generate an answer to a query which follows the NPOV principle. We quote here an extract of Wikipedia’s NPOV definition [^2] The entire definition is very informative, and we invite the readers to look through Appendix A for more details. All of our writers read the entire definition, and attempt to follow when developing their examples. :
"convey to the reader the information […] fairly, proportionately, and as far as possible without editorial bias. Wikipedia aims to describe disputes, but not engage in them."
One fundamental goal of this work is to combat polarization by developing a deeper understanding of different points of view and empathy for the people who express them. As such, we strive to ensure that people’s views are represented; this can be seen as a form of reflective listening, which has been shown to decrease defensiveness and increase willingness to hear different perspectives Dalmar (2007). As a consequence, we do not assume a hierarchy between scientific arguments and emotional or spiritual arguments; this is in stark contrast with the goal of Wikipedia or any encyclopedia. In our definition of NPOV, by design, peer-reviewed quantitative studies can be presented on equal ground with widespread moral positions.
@@ 3.2 Defining the NPOV task
3.3 Methodology for writing examples
We spent significant effort on the methodology to create high quality reward model training examples (achieving high inter-annotator agreement with 90% of annotations within distance 1 from each other on a 0-5 scale), and a reward model with 99.3% AUC), specifically we:
  • Gather a diverse team, to help uncover and mitigate writers’ biases.
  • Choose controversial topics. The SHQ-NPOV dataset focuses on US-centric topics.
  • Research the topics and cross-reference multiple sources.
  • Write answers, peer-critique them and revise based on critics until agreement.
  • Drop examples if no agreement can be found after 2 rounds.
  • Occasionally write non-NPOV answers, which are also useful to train reward-based methods.
See Appendix D for a detailed description of our methodology.
3.4 Crowd-annotation procedure
@@
To validate our generated answers, we hired a team of 10 external crowd-annotators based in the US and recruited 4 internal experts, for a total of 14 annotators. 2 experts first annotated the whole dataset, using a score from 0 to 5 to capture "NPOVness" (see Table 5 for score meaning). After reading the Wikipedia definition for NPOV, crowd-annotators are then split into 2 groups, and trained for the task on half the dataset. They are then asked to rate the NPOVness of an answer for the other half. At least 3 annotators rate each query-answer pair across all experiments, for a maximum of 5. The final NPOV score is the average of the 3–5 annotators’ scores. See Appendix K for more details.
Figure 2: Distribution of L1 distance between annotators’ score and NPOV score. 75% of answers have less than 0.50 difference with the NPOV score, and 90% of answers less than 1.
@@
The annotators have high agreement, as seen in the metrics below. When we compute a metric with respect to the NPOV score (defined as the average of the annotator’s score), the metric is unfairly high, as every annotator contributes to this score. Similarly, when we compute a metric with respect to the score of a single expert, the metric is unfairly low. We can therefore make the following claims:
  • The overall annotator average accuracy is within $91.07\%,91.15\%]$.
  • The overall annotator average AUC is within $96.53\%,97.59\%]$.
@@
The last point is of particular interest. Most annotators’ score is less than 0.5 away from the real score, which is very precise for a 0–5 score. We plot the exact distribution of L1 distances in Figure 2.
3.5 1190 extra queries for evaluation
The SHQ-NPOV dataset consists of 300 query-answer pairs covering 89 topics, which we call in-distribution topics (id topics). Additionally, we also build a query dataset for evaluation using both these topics and 30 out-of-distribution ones (ood topics), making a total of 119 that we upsample using the 10 templates from Table 6, giving a total of 1190 evaluation queries. Note that there is no overlap between these additional queries and the queries in the SHQ-NPOV dataset.
@@
4 Experimental setup
4.1 Methods and baselines
We now describe the different methods we compare for generating NPOV answers. Note that all the methods are used to tune the same base model (about 20B parameters). The budget for all the experiments was about of 50k TPUs hours. We compare the results of PE-RL (Parameter-Efficient Reinforcement Learning) to the following baselines: base model, LoRA (parameter-efficient Supervised Fine Tuning), SFT ("traditional" Supervised Fine Tuning), RLHF ("traditional" Reinforcement Learning with Human Feedback) and best-of-10. For each of these techniques, we also consider its variant "+ preamble", which consists of adding the preamble in Appendix E.1 to the prompt before the query.
PE-RL:
Following the method introduced in Sidahmed et al. (2024), we train 2 LoRA adapters, one for the reward model and one for the policy. We first train the reward model by finetuning LoRA Hu et al. (2022) on the  20B base model. We use a rank-4 LoRA adapter (about 4M trainable parameters), a learning rate of $0.0005$ for $8,000$ steps to train on all the 300 rated examples. We then optimize the RL loop and update the parameters of the rank-4 LoRA adapter for the policy using the trained reward model for $2,000$ steps (with 1,000 warm-up steps), using a value learning rate of $0.0001$, policy learning rate of $0.00001$. We perform a hyper-parameters search to pick the best model. The LoRA adapter for the policy is initialized at 0 at the begininng of the RL loop.
PE-RL + LoRA SFT:
Instead of initializing the LoRA adapter for the policy with a null checkpoint, we perform LoRA SFT on the 132 "great" (query, answer) pairs, and initialize the RL training with this trained checkpoint. Note that the 132 examples are part of the 300 used for training the reward model, so this method does not use additional data.
LLM $\sim$20B :
We report the generations from the base model, an instruction-tuned LLM of about 20B parameters.
LoRA SFT:
We fine-tuned the base model with a rank-4 LoRA Hu et al. (2022) adapter (about 6M trainable parameters) using only examples rated as sufficient or above ($N=132$), a batch size of $64$, a learning rate of $0.0001$, and dropout probability of $0.1$ for $800$ steps. The set of hyper-parameters is selected from a grid search of LoRA ranks in $\{1,4,8,16\}$, learning rate in $\{0.0004,0.005,0.001,0.05\}$, and dropout probability in $\{0,0.05,0.1,0.2\}$.
SFT:
We fine-tuned the base LLM exclusively on examples with ratings of sufficient or higher ($N=132$). The model was trained for $1000$ steps using a batch size of $64$, a learning rate of $0.0001$, and a dropout rate of $0.1$. These hyper-parameters were chosen based on a grid search over learning rates $\{0.0001,0.001,0.005,0.05\}$ and dropout rates $\{0,0.05,0.1,0.2\}$.
RLHF: